### v1

# Optimal energy storage and collective charging speedup in the central-spin quantum battery

Abstract
Quantum batteries (QBs) exploit principles of quantum mechanics to accelerate the charging process and aim to achieve optimal energy storage. However, analytical results for investigating these problems remain lacking due to the challenges associated with nonequilibrium dynamics. In this work, we analytically investigate a central-spin QB model in which $N_{b}$ spin-1/2 battery cells interact with $N_{c}$ spin-1/2 charger units, using $m$ initially excited charger units as a resource. By employing the invariant subspace method and the shifted Holstein-Primakoff (HP) transformation, we identify four scenarios in which optimal energy storage can be achieved: 1) $N_{b}\!\ll\!m\!\ll\!N_{c}$; 2) $m\!\ll\!N_{b}\!\ll\!N_{c}$; 3) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and 4) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$]. In these cases, optimal storage is ensured by the SU(2) symmetry emerging from the charging dynamics. The first three cases map the central-spin QB to different Tavis-Cummings (TC) QBs, while the fourth corresponds to the non-TC limit. We analytically determine the charging time and demonstrate that in the fully charging cases (1) and (4), the collective charging exhibits an $N_{b}$-fold enhancement in speedup compared to the parallel charging scheme. Additionally, we numerically observe a unified charging behavior when $m\!=\!N_{c}$, showing that asymptotically optimal energy storage is possible when $N_{b}\!=\!m\!=\!N_{c}$. In this case, we find a collective charging enhancement scaling as $N_{b}^{0.8264}$. Our results highlight the crucial role of dynamically emergent SU(2) symmetry in providing an analytical understanding of non-equilibrium charging dynamics in QBs.
Figure 1: Schematic diagram of the central-spin QB. (a) Initially, the system consists of $N_{b}$ battery cells in their ground states, with no excitations, and $N_{c}$ charger units prepared in a Dicke state containing $m$ excitations. Comparison between (b) the collective charging scheme and (c) the parallel charging scheme.
I Introduction
Pioneering theoretical work on energy extraction from many-body states has suggested that the generation of multipartite entanglement can enhance the efficiency of work extraction Hovhannisyan et al. (2013). This finding implies the existence of a new class of energy storage devices, known as quantum batteries Alicki and Fannes (2013) (QBs), which exploit unique quantum features to accelerate the charging process and outperform classical batteries. Compared to charging individual battery cells independently, collective charging exhibits a significant quadratic speedup in charging time Gyhm et al. (2022); Campaioli et al. (2017); Julià-Farré et al. (2020), which is believed to arise from multipartite entanglement induced by collective interactions. The phenomenon where charging power scales faster than the number of quantum cells is referred to as collective charging advantage Campaioli et al. (2024), which has been observed in various models, including central-spin model Peng et al. (2021), spin-chain model Le et al. (2018); Huangfu and Jing (2021), Sachdev–Ye–Kitaev model Rossini et al. (2020) and Dicke model Zhang and Blaauboer (2023); Ferraro et al. (2018); Dou et al. (2022).
On the other hand, energy storage is another crucial metric for evaluating the performance of a quantum battery. Practically, a QB that charges rapidly but stores only minimal energy is not desirable. Recent studies have explored the relationship between energy storage and quantum correlations Shi et al. (2022); Andolina et al. (2019); Šafránek et al. (2023). However, identifying many-body models that satisfy the conditions for achieving optimal energy storage remains challenging. Here, optimal energy storage refers to a situation where the battery absorbs all the energy initially stored in the charger or reaches a fully charged state, which will be rigorously defined later. Our recent work has shown that the emergence of SU(2) symmetry in the Tavis-Cummings (TC) model guarantees the realization of optimal energy storage Yang et al. (2024). Could this emergent symmetry also explain the presence of optimal energy storage in other models? Moreover, under conditions of optimal energy storage, can we achieve collective charging speedup simultaneously? We aim to address these questions in an important low-dimensional model.
The central-spin model has provided significant insights into the quantitative understanding of decoherence in nitrogen-vacancy centers in diamond, as well as the dynamics of entanglement Schliemann et al. (2003); Quan et al. (2006); Hanson et al. (2008); Bortz et al. (2010); Barnes et al. (2012); Lu et al. (2020); Wan et al. (2020); Claeys et al. (2018); Wu et al. (2020); He et al. (2019); Stanek et al. (2013); Cucchietti et al. (2005); Lee et al. (2008); Onizhuk et al. (2024); Rossini et al. (2007). This model can be naturally put into the framework of QBs, referred to as the central-spin QB, where $N_{b}$ central spins act as battery cells and $N_{c}$ bath spins serve as charging units. It has been argued that a collective charging advantage is present in this model; however, the analysis is based on applying the Holstein-Primakoff (HP) transformation to both the bath and central spins Peng et al. (2021), which is valid only in the short-time regime and does not extend to the time scale including maximal energy storage Andolina et al. (2019). On the other hand, the use of the HP approximation simplifies the central-spin model to the TC model, which we refer to as the TC limit. As discussed in Ref. Yang et al. (2024), even within the TC model, optimal energy storage is not guaranteed in all cases. In a specific non-TC limit Liu et al. (2021), we identified an optimal energy storage scenario for the case of $N_{b}\!=\!2$ and conjectured, based on numerical evidence, that optimal energy storage persists under this condition for arbitrary $N_{b}$ case Liu et al. (2021). The corresponding charging time has also been conjectured and numerically validated. Despite these results, analytical results regarding the conditions for optimal energy storage and the determination of charging time in the central-spin model remain lacking. In particular, can we provide rigorous proof for the conditions and charging time for optimal energy storage in the non-TC limit? Additionally, even for the TC limit, different initial configurations in the central-spin model can be used to realize it. Do these various configurations all guarantee optimal energy storage? Do they yield the same charging time? Addressing these questions will help distinguish the central-spin model from its behavior under the TC limit and deepen our understanding of the nonequilibrium dynamics in the central-spin model.
In this paper, we address these questions systematically. In Sec. II, we introduce the central-spin model and the invariant subspace approach used for the subsequent analysis. The quantities that characterize optimal energy storage and the collective charging advantage are also given. Section III employs the HP transformation to identify four initial configurations that achieve the TC limits. For the first three cases, we rigorously demonstrate the emergent SU(2) symmetry that ensures optimal energy storage and analytically determine the corresponding charging time. Based on these findings, we prove that the collective charging advantage scales with $N_{b}$ in the first fully charged TC limit. For the fourth case, numerical calculations indicate the absence of optimal energy storage. In Sec. IV, we explore the non-TC limits. We analytically demonstrate a novel scenario for achieving optimal energy storage while maintaining an $N_{b}$ scaling for the collective charging advantage. Additionally, numerical analysis reveals another non-TC configuration that attains asymptotic optimal energy storage, with a collective charging advantage scaling as $N_{b}^{0.8264}$. Finally, a conclusion is made in Sec. V.
II Central-spin quantum battery
The central-spin QB is described by the following Hamiltonian
$H$ $=$ $H_{b}+H_{c}+H_{I},$
$H_{b}$ $=$ $\omega_{b}S^{z},$
$H_{c}$ $=$ $\omega_{c}J^{z},$
$H_{I}$ $=$ $A(S^{+}J^{-}+S^{-}J^{+}),$ (1)
where $H_{b}$, $H_{c}$, and $H_{I}$ correspond to the Hamiltonian of the battery, charger, and their homogeneous flip-flop spin interactions in the $x$-$y$ plane, respectively. Here, $S^{\alpha}\!=\!\sum_{j=1}^{N_{b}}\sigma_{j}^{\alpha}/2$ and $J^{\alpha}\!=\!\sum_{k=1}^{N_{c}}\sigma_{k}^{\alpha}/2$ with $\alpha\!=\!x,y,z$ are the collective spin operators for the $N_{b}$ spin-1/2 battery cells and $N_{c}$ spin-1/2 charger units, respectively. The spin ladder operators are given by $J^{\pm}\!=\!J^{x}\pm iJ^{y}$ and $S^{\pm}\!=\!S^{x}\pm iS^{y}$.
Initially, the battery is set to the ground state of $H_{b}$ with $N_{b}$ spins down, namely, $|0\rangle_{b}\!\equiv|\!\downarrow_{1},\downarrow_{2},\cdot\cdot\cdot,\downarrow_{N_{b}}\rangle_{b}$, while the charger is prepared in a high-energy Dicke state $|m\rangle_{c}$ with $m$ spins up, see Fig.1(a). The charging dynamics are performed by turning on the interactions $H_{I}$. To ensure that the energy injected into the battery exclusively comes from the charger, we require $[H_{b}+H_{c},H]\!=\!0$ and thus we only consider the resonant case with $\omega_{b}\!=\!\omega_{c}\!\equiv\!\omega$. Therefore, the free part in the Hamiltonian (II) is a constant and we will ignore it. In terms of the invariant subspace $\mathcal{H}_{m}=\{|0\rangle_{b}|m\rangle_{c},|1\rangle_{b}|m\!-\!1\rangle_{c},\cdot\cdot\cdot,|d\rangle_{b}|m\!-\!d\rangle_{c}\}$ containing the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$, the Hamiltonian $H$ defined in Eq. (II) can be represented as a $(d+1)\times(d+1)$ matrix in the invariant subspace $\mathcal{H}_{m}$ by using Dicke basis
$\bm{H}=\begin{pmatrix}0&u_{1}&\\ u_{1}&0&u_{2}&\\ &\ddots&\ddots&\ddots&\\ &&u_{d-1}&0&u_{d}\\ &&&u_{d}&0\end{pmatrix},$ (2)
where we have shifted the ground state energy to cancel the diagonal terms. The off-diagonal terms are given by
$u_{j}(N_{b},N_{c},m)$
$=A\sqrt{j(N_{b}-j+1)(N_{c}-m+j)(m-j+1)}.$ (3)
The dimension of the effective Hamiltonian (2) is $(d\!+\!1)$ with $d\equiv\min\{{N_{b},m}\}$, which scales linearly with the minimal number of battery cells or the charger units. Therefore, it is easy to diagonalize $\bm{H}$ for calculating dynamics $|\psi(t)\rangle=e^{-iHt}|\psi(0)\rangle$, details can be found in Ref. Liu et al. (2021).
The energy transferred from the charger to the battery is given by
$\Delta E(t)=\langle\psi(t)|H_{b}|\psi(t)\rangle-\langle\psi(0)|H_{b}|\psi(0)\rangle,$ (4)
[…]
$\eta(t)\equiv\frac{\Delta E(t)}{\langle\psi_{\rm opt}|H_{b}|\psi_{\rm opt}\rangle-\langle\psi(0)|H_{b}|\psi(0)\rangle},$ (5)
where the denominator represents the optimal energy storage, given by $\omega\min\{N_{b},m\}$, and $|\psi_{\rm opt}\rangle\!=\!|\min\{N_{b},m\}\rangle_{b}|m\!-\!\min\{N_{b},m\}\rangle_{c}$ is the optimal charged state. Therefore, the condition for optimal charging can be expressed as $\eta(T)\!=\!1$. One of our main goals is to analytically determine which types of initial states can achieve optimal energy storage.
Under the restriction of optimal energy storage, our second goal is to demonstrate the collective charging advantage analytically. In the collective charging scheme, shown in Fig.1(b), $N_{b}$ battery cells collectively interact with $N_{c}$ charging units, with the charger initially having $m$ spins in the excited state. The collective charging power is defined as
$P_{\#}(N_{b},N_{c},m)\equiv\frac{\Delta E(T)}{T(N_{b},N_{c},m)},$ (6)
which depends solely on the battery size $N_{b}$, the charger size $N_{c}$, and the initial number of excited charger spins $m$. For a fair comparison, the resources used in the parallel charging scheme [see Fig. 1(c)] must be equivalent to those used in the collective charging scheme. In the parallel charging scheme, there are $N_{b}$ individual single-cell QBs, where each battery cell interacts with $N_{c}/N_{b}$ charger units, with $m/N_{b}$ spins initially excited. To make sense for this setting, we require that $N_{c}\!\geq\!N_{b}$ and $m\!\geq\!N_{b}$, and we will define the collective charging advantage Eq. (9) only for this case. According to Ref. Liu et al. (2021), the charging power for a single-cell central-spin QB is given by
$P_{\rm single\!-\!cell}$ $=$ $\frac{2\omega}{\pi}u_{1}(1,N_{c}/N_{b},m/N_{b})$ (7)
$=$ $\frac{2\omega A}{\pi}\sqrt{\frac{m}{N_{b}}\left(\frac{N_{c}}{N_{b}}-\frac{m}{N_{b}}+1\right)}.$
In the parallel charging scheme, the charging time remains the same as in the single-cell QB case, but the total stored energy will be $N_{b}$ times larger. Therefore, the parallel charging power is given by
$P_{\parallel}\equiv N_{b}P_{\rm single\!-\!cell}.$ (8)
The collective charging advantage is thus quantified by
$\Gamma=\frac{P_{\#}}{P_{\parallel}}=\frac{P_{\#}}{N_{b}P_{\rm single\!-\!cell}}.$ (9)
If $\Gamma\!\sim\!N_{b}^{\alpha}$ with $\alpha\!>\!0$, we say the collective charging advantage exists.
Figure 2: Charging dynamics for four TC limits: a) $N_{b}\!\ll\!m\!\ll\!N_{c}$; b) $m\!\ll\!N_{b}\!\ll\!N_{c}$; c) $m\!\ll\!N_{c}\!\ll\!N_{b}$; and d) $m\!=\!N_{c}\!\ll\!N_{b}$. The blue lines are numerically obtained and the red lines are theoretical results. The coupling strength is set to $A\!=\!1$.
III Optimal energy storage and collective charging speedup at TC Limits
In previous work Liu et al. (2021), we analytically demonstrated a case of optimal energy storage for $2\!=\!N_{b}\!\ll\!m\!\ll\!N_{c}$ and numerically verified the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$ with arbitrary $N_{b}$ to achieve optimal energy storage. However, the corresponding charging time was not determined. Actually, under these initial conditions, the central-spin model can be mapped to the TC model, which we refer to as the TC limit. In this section, we will extend our previous work to find more TC limits and then we will use the theory developed in Ref. Yang et al. (2024) to determine not only the possibility for achieving optimal energy storage but also the collective charging advantage. We emphasize that the emergent SU(2) symmetry in the dynamics plays a crucial role in achieving optimal energy storage and is key to analytically calculating the collective charging advantage.
For the case of $N_{b}\!\ll\!m\!\ll\!N_{c}$, due to the conservation of the total excitation number $[J^{z}\!+\!S^{z},H]\!=\!0$, the number of excitations in the charger part cannot exceed the initial value $m$. Since $\lim_{N_{c}\to\infty}m/N_{c}=0$, the charger remains close to the vacuum state $|0\rangle_{c}$ throughout the dynamics at the limit $N_{c}\!\to\!\infty$ Therefore, we can apply the HP transformation to the charger Hamiltonian as follows
$J^{+}\to\sqrt{N_{c}}a^{\dagger}\sqrt{1-\frac{a^{\dagger}a}{N_{c}}},$
$J^{z}\to-\frac{N_{c}}{2}+a^{\dagger}a,$ (10)
where $\sqrt{1-a^{\dagger}a/N_{c}}\simeq 1$ because $\langle\psi(t)|a^{\dagger}a|\psi(t)\rangle/N_{c}\!=\!\langle\psi(t)|J^{z}|\psi(t)\rangle/N_{c}\!+\!1/2\!\leq\!m/N_{c}\!\to\!0$. Under this approximation, the original Hamiltonian (II) reduces to
$H\overset{N_{b}\!\ll\!m\!\ll\!N_{c}}{\longrightarrow}A\sqrt{N_{c}}(S^{+}a+S^{-}a^{\dagger}),$ (11)
where $a$ and $a^{\dagger}$ are bosonic annihilation and creation operators. In this correspondence, the number of charger spin excitations $m$ in the initial state $|\psi(0)\rangle\!=\!|0\rangle_{b}|m\rangle_{c}$ corresponds to the photon number. In Ref. Yang et al. (2024), we have shown that for TC battery (11) and initial state $|\psi_{\rm TC}(0)\rangle=|0\rangle_{b}|n_{0}\rangle_{c}$ with $|n_{0}\rangle$ being a Fock state with $n_{0}$ photons, a distinct SU(2) symmetry emerges from the dynamics for the limiting cases $n_{0}\!\gg\!N_{b}$ or $N_{b}\!\gg\!n_{0}$. This results in optimal energy storage given by $\Delta E(T_{\rm TC})\!=\!\omega\min\{N_{b},n_{0}\}$ with the charging time
$T_{\rm TC}=\frac{\pi}{2A\sqrt{N_{c}}\sqrt{\max\{N_{b},n_{0}\}-\frac{\min\{N_{b},n_{0}\}-1}{2}}}.$ (12)
Therefore, we conclude that optimal energy storage in the central-spin QB can be achieved with the initial state $|\psi(0)\rangle=|0\rangle_{b}|m\rangle_{c}$ with $N_{b}\!\ll\!m\!\ll\!N_{c}$, see Fig.2(a). By substituting $n_{0}\!=\!m$ into Eq. (12), the charging time is given by
$T(N_{b}\!\ll\!m\!\ll\!N_{c})=\frac{\pi}{2A\sqrt{N_{c}}\sqrt{m-\frac{N_{b}-1}{2}}},$ (13)
which is numerically verified in Fig. 2(a). Then by substituting Eqs. (7) and (13) into Eq. (9), the collective charging advantage is obtained as follows
$\Gamma(N_{b}\!\ll\!m\!\ll\!N_{c})=\frac{\sqrt{N_{c}}\sqrt{m-\frac{N_{b}-1}{2}}}{\sqrt{\frac{m}{N_{b}}\left(\frac{N_{c}}{N_{b}}-\frac{m}{N_{b}}+1\right)}}\simeq N_{b},$ (14)
which indicates an $N_{b}$-fold collective charging speedup compared to the parallel charging scheme.
It is important to emphasize that the TC Hamiltonian (11), which describes the central-spin QB, only requires the number of excited charger spins to be much smaller than the total number of charger spins, i.e., $m\!\ll\!N_{c}$. Moreover, the condition $n_{0}\!\ll\!N_{b}$ for optimal energy storage in the TC battery implies another scenario: $m\!\ll\!N_{b}\!\ll\!N_{c}$, which also facilitates optimal energy storage in the central-spin QB, see Fig. 2(b). By setting $n_{0}\!=\!m$ in Eq. (12), we obtain the charging time as follows
$T(m\!\ll\!N_{b}\!\ll\!N_{c})=\frac{\pi}{2A\sqrt{N_{c}}\sqrt{N_{b}-\frac{m-1}{2}}},$ (15)
which is numerically verified in Fig. 2(b). Note that calculating the charging advantage in this case makes no sense because $m/N_{b}\!\to\!0$. Therefore, we will only consider the charging advantage for $N_{b}\!\leq\!m$ and $N_{b}\!\leq\!N_{c}$ in the following discussion.
Unlike the two situations discussed above, we can also employ the HP transformation to map the battery onto a bosonic mode, facilitating the construction of a TC QB from the central-spin QB. Now, let us consider the case where $m\!\ll\!N_{c}\!\ll\!N_{b}$. In this regime, the possible maximum excitation for the battery part is $m$, which is significantly smaller than the battery size $N_{b}$. Consequently, the HP transformation can be applied to the battery spins, yielding
$S^{+}\to\sqrt{N_{b}}b^{\dagger}\sqrt{1-\frac{b^{\dagger}b}{N_{b}}},$
$S^{z}\to-\frac{N_{b}}{2}+b^{\dagger}b,$ (16)
where, by approximating $\sqrt{1-b^{\dagger}b/N_{b}}\simeq 1$, the central-spin QB in Eq. (II) reduces to the TC QB,
$H\overset{m\!\ll\!N_{c}\!\ll\!N_{b}}{\longrightarrow}A\sqrt{N_{b}}(bJ^{+}+b^{\dagger}J^{-}),$ (17)
where the battery has become a photon field. From the perspective of matrix representation in the invariant subspace, Eq. (2), the above HP transformation approximates the matrix element $u_{j}$ as
$u_{j}$ $=$ $A\sqrt{j(N_{b}-j+1)(N_{c}-m+j)(m-j+1)}$ (18)
$\overset{m\!\ll\!N_{b}}{\longrightarrow}$ $A\sqrt{N_{b}}\sqrt{j(N_{c}-m+j)(m-j+1)}$
with $j\!=\!1,\ldots,m$. Considering the additional condition $m\ll N_{c}$, we can further approximate Eq. (18) as
$u_{j}$ $\overset{m\!\ll\!N_{c}\!\ll\!N_{b}}{\longrightarrow}$ $A\sqrt{N_{b}}\sqrt{N_{c}-\frac{m-1}{2}}\sqrt{j(m-j+1)},$ (19)
which implies that the effective Hamiltonian (2) is exactly the generator of the SU(2) algebra
$\bm{H}\overset{m\!\ll\!N_{c}\!\ll\!N_{b}}{\longrightarrow}\sqrt{N_{b}}\Omega\bm{L^{x}},$ (20)
where $\Omega\equiv 2A\sqrt{N_{c}-(m-1)/2}$ is the generalized Rabi frequency and $\bm{L^{x}}$ is $x$-direction generator of SU(2) algebra in the spin-$m$ representation. The evolution of energy storage (4) can be represented as
$\Delta E(t)=\frac{m\omega}{2}+\bm{\psi}^{\dagger}(0)\bm{L^{z}}(t)\bm{\psi}(0)\omega,$ (21)
where $\bm{L^{z}}(t)\!=\!\exp(it\Omega\sqrt{N_{b}}\bm{L^{x}})\bm{L^{z}}\exp(-it\Omega\sqrt{N_{b}}\bm{L^{x}})$ and $\bm{\psi}(0)\!=\!(1,0,\cdots,0)^{T}$. Using the Baker-Campbell-Hausdorff formula, we have
$\bm{L^{z}}(t)=\sin(\Omega\sqrt{N_{b}}t)\bm{L^{y}}+\cos(\Omega\sqrt{N_{b}}t)\bm{L^{z}},$ (22)
and thus
$\Delta E(t)=\frac{m\omega}{2}[1-\cos(\Omega\sqrt{N_{b}}t)],$ (23)
which implies that optimal energy storage can be achieved at the charging time
$T(m\!\ll\!N_{c}\!\ll\!N_{b})=\frac{\pi}{\Omega\sqrt{N_{b}}}$
$=\frac{\pi}{2A\sqrt{N_{b}}\sqrt{N_{c}-\frac{m-1}{2}}},$ (24)
see Fig.2(c) for numerical verification.
However, we claim that optimal energy storage cannot be achieved in all TC limits, where the emergence of SU(2) symmetry plays a crucial role. As a specific case, we consider the scenario where $m\!=\!N_{c}\!\ll\!N_{b}$. Given the ratio $m/N_{b}\to 0$, the maximum possible number of excitations in the battery during the dynamics, i.e., $m$, is close to zero. Consequently, the effective Hamiltonian remains identical to Eq.(17). Nevertheless, since $m=N_{c}$ we are unable to apply the same approximation used in deriving Eq. (19). As a result, it is not possible to obtain an SU(2) generator to describe the charging dynamics in this particular TC limit. The absence of SU(2) symmetry in the regime $m\!=\!N_{c}\!\ll\!N_{b}$ therefore prevents the realization of optimal energy storage, as shown in Fig. 2(d).
Figure 3: (a) Charging dynamics in the non-TC limit: $N_{b}\!\ll\!m\!=\!kN_{c}$ with $10\!<\!k\!<\!$ 1. The blue lines represent numerical results, while the red lines correspond to theoretical predictions. (b) The analytical charging time, Eq. (33), as a function of $k$ is numerically verified. The coupling strength is set to $A\!=\!1$.
IV Optimal energy storage and collective charging speedup at non-TC Limits
Having successfully applied the SU(2) symmetry theory to address the optimal energy storage problem in the TC limits, in this section, we will use it again to investigate the problem in the non-TC limits. The non-TC limits distinguish the central-spin QB from the TC QB, representing a unique regime specific to the central-spin QB.
The first non-TC limit considered here is $N_{b}\!\ll\!m\!=\!kN_{c}$ with $k\!\in\!(0,1)$. Since $N_{b}\ll m$, the charger part can be considered essentially unchanged during the dynamics. Therefore, macroscopic excitations are always present in the charger, i.e.,
$\lim_{N_{c}\to\infty}\frac{\langle\psi(t)|\hat{a}^{\dagger}a|\psi(t)\rangle}{N_{c}}$
$=\lim_{N_{c}\to\infty}\frac{\langle\psi(t)|J^{z}|\psi(t)\rangle}{N_{c}}+\frac{1}{2}=\frac{m}{N_{c}}=k,$ (25)
which prevents the direct application of the HP transformation given in Eq. (III). To circumvent this issue, we can first apply the HP transformation, Eq. (III), and then shift the bosonic operators as follows
$\hat{a}=\hat{c}+\sqrt{kN_{c}}.$ (26)
With this shift, we obtain
$\lim_{N_{c}\to\infty}\frac{\langle\psi(t)|\hat{c}^{\dagger}c|\psi(t)\rangle}{N_{c}}$
$=\lim_{N_{c}\to\infty}\frac{\langle\psi(t)|\hat{a}^{\dagger}a|\psi(t)\rangle}{N_{c}}-\frac{kN_{c}}{N_{c}}=0.$ (27)
Figure 4: (a) The maximum energy storage efficiency $\eta(T)$ as a function of $N_{c}/N_{b}$ for the case $m\!=\!N_{b}$. (b) Charging dynamics for $m\!=\!N_{b}\!=\!N_{c}=100$. The coupling strength is set to $A\!=\!1$.
Substituting Eq. (26) into Eq. (III), we have
$J^{+}$ $=$ $\sqrt{(1-k)N_{c}}(c^{\dagger}+\sqrt{kN_{c}})$ (28)
×$\times\sqrt{1-\frac{c^{\dagger}c+\sqrt{kN_{c}}(c+c^{\dagger})}{(1-k)N_{c}}}$
$=$ $\sqrt{k(1-k)}N_{c}+\mathcal{O}(\sqrt{N_{c}}),$
where we used Eq. (IV). To leading order in $N_{c}$, the central-spin model in Eq. (II) reduces to
$H$ $\overset{N_{b}\!\ll\!m=kN_{c}}{\simeq}$ $A\sqrt{k(1-k)}N_{c}(S^{-}+S^{+})$ (29)
$=$ $2A\sqrt{k(1-k)}N_{c}S^{x},$
which is exactly the generator of the SU(2) algebra, and therefore means that SU(2) symmetry also appears in this non-TC limit. The evolution of energy storage (4) in this case can be expressed as
$\Delta E(t)=\frac{N_{b}\omega}{2}+\langle\psi(0)|S^{z}(t)|\psi(0)\rangle,$ (30)
where $S^{z}(t)\!=\!\exp(itH)S^{z}\exp(-itH)$. Using the Baker-Campbell-Hausdorff formula, we find
$S^{z}(t)$ $=$ $\sin(2A\sqrt{k(1-k)}N_{c}t)S^{y}$ (31)
$+\cos(2A\sqrt{k(1-k)}N_{c}t)S^{z}.$
Thus, the energy evolution is given by
$\Delta E(t)=\frac{N_{b}\omega}{2}[1-\cos(2A\sqrt{k(1-k)}N_{c}t)],$ (32)
which indicates that optimal energy storage is achieved at the charging time
$T(N_{b}\!\ll\!m\!=\!kN_{c})=\frac{\pi}{2AN_{c}\sqrt{k(1-k)}},$ (33)
as confirmed numerically in Fig.3(b). Verification of optimal energy storage is shown in Fig. 3(a). By substituting Eq. (33) into Eq. (9), we obtain the collective charging advantage as
$\Gamma(N_{b}\!\ll\!m\!=\!kN_{c})=\frac{N_{c}\sqrt{k(1-k)}}{\sqrt{\frac{m}{N_{b}}\left(\frac{N_{c}}{N_{b}}-\frac{m}{N_{b}}+1\right)}}\simeq N_{b}.$ (34)
In summary, we show that SU(2) symmetry can also emerge in the non-TC limit, leading to optimal energy storage with a collective charging advantage scaling as $N_{b}$.
Figure 5: Scaling of the collective charging advantage $\Gamma$ as a function of the number of battery cells $N_{b}$ in the case of $N_{b}\!=\!m\!=\!N_{c}$. The coupling strength is set to $A\!=\!1$.
Note that Eq. (33) does not hold for the case $k\!=\!1$, or equivalently $m\!=\!N_{c}$, because the approximation of $J^{+}$ is incorrect in this scenario, as the zeroth-order term vanishes. To investigate the region where $m\!=\!N_{c}$, we perform numerical calculations, see Fig. (4). In Fig. 4(a) we obtain the maximum energy transferred into the battery and plot the energy storage efficiency $\eta(T)$, defined in Eq. (5), at the charging time $T$ for various values of $N_{b}$. Surprisingly, $\eta(T)$ exhibits a universal behavior, with all curves collapsing onto a single curve after rescaling. Moreover, we observe that $\eta(T)\!\to\!1$ for the case $N_{c}\!=\!N_{b}$, indicating the possibility of achieving optimal energy storage. In Fig. 4(b), we plot the charging dynamics for $N_{b}\!=\!m\!=\!N_{c}$ which shows asymptotic optimal energy storage, meaning that the first maximum of $\eta$ approaches one. However, due to the absence of dynamically emergent SU(2) symmetry, this case differs significantly from the previous situations and does not exhibit sine-like oscillations. In Fig. 5, we plot the collective charging advantage for $N_{b}\!=\!m\!=\!N_{c}$ and find that $\Gamma\propto N_{b}^{0.8264}$. Although $\Gamma$ does not scale exactly with $N_{b}$ in this case, it still demonstrates a collective charging advantage close to an $N_{b}$-scaling behavior.
Table 1: Optimal energy storage, collective charging advantage, and emergent SU(2) symmetry in the central-spin QB.
Cases Charging Time Optimal Energy Storage Emergence SU(2) Symmetry Collective Charging Advantage
$N_{b}\!\ll\!m\!\ll\!N_{c}$ $\frac{\pi}{2A\sqrt{N_{c}}\sqrt{m-\frac{N_{b}-1}{2}}}$ Yes Yes $N_{b}$
$m\!\ll\!N_{b}\!\ll\!N_{c}$ $\frac{\pi}{2A\sqrt{N_{c}}\sqrt{N_{b}-\frac{m-1}{2}}}$ Yes Yes NA
$m\!\ll\!N_{c}\!\ll\!N_{b}$ $\frac{\pi}{2A\sqrt{N_{b}}\sqrt{N_{c}-\frac{m-1}{2}}}$ Yes Yes NA
$m\!=\!N_{c}\!\ll\!N_{b}$ Numerical No No NA
$N_{b}\!\ll\!m\!=\!kN_{c}$
$k\!\in\!(0,1)$ $\frac{\pi}{2AN_{c}\sqrt{k(1-k)}}$ Yes Yes $N_{b}$
$m\!=\!N_{c}\!=\!N_{b}$ Numerical Asymptotically No $N_{b}^{0.8264}$
V Conclusion
In this work, we investigate the charging dynamics of a central-spin QB. The problems of optimal energy storage and collective charging advantage are analytically explored within this model. We rigorously demonstrate the realization of optimal energy storage in four different initial state settings: 1) $N_{b}\!\ll\!m\!\ll\!N_{c}$; 2) $m\!\ll\!N_{b}\!\ll\!N_{c}$; 3) $m\!\ll\!N_{c}\!\ll\!N_{b}$; 4) $N_{b}\!\ll\!m\!=\!kN_{c}$ [$k\!\in\!(0,1)$], where $N_{b}$ denotes the number of battery cells, $N_{c}$ represents the number of charger units, and $m$ is the initial number of excited charger units. The first three cases correspond to the TC limits, while the fourth case represents the non-TC limit. Our findings reveal that, across all these cases, the emergence of SU(2) symmetry from the charging dynamics is crucial for achieving optimal energy storage. This emergent symmetry also provides a theoretical approach to analytically determine the charging time. In the fully charged scenarios corresponding to cases (1) and (4), a collective charging speedup of $N_{b}$-fold is proven in comparison to the parallel charging scheme. Furthermore, in the case of $N_{b}\!=\!m\!=\!N_{c}$ we identify the possibility of asymptotically achieving optimal energy storage, with numerical calculations indicating a collective charging advantage scaling as approximately $\sim N_{b}^{0.8264}$. The main results are summarized in Table 1.
Our work highlights the significance of emergent SU(2) symmetry in analytically addressing the challenges of optimal energy storage and collective charging speedup. It extends previous studies on the TC QB Yang et al. (2024) and the central-spin QB Liu et al. (2021). Especially in cases (1) and (4), not only the optimal energy storage can be achieved, but also the collective charging advantage of $N_{b}$ exists, which provides guidance for designing high-performance QBs. We believe the methods used here can be applied to other QB models to uncover further the mechanisms underlying optimal energy storage and collective charging speedup.
ACKNOWLEDGMENTS
This work was supported by the NSFC (Grants No. 12275215, No. 12305028, and No. 12247103), Shaanxi Fundamental Science Research Project for Mathematics and Physics (Grant No. 22JSZ005) and the Youth Innovation Team of Shaanxi Universities. HLS was supported by the European Commission through the H2020 QuantERA ERA-NET Cofund in Quantum Technologies project “MENTA”.